# Exploiting the Semantic Knowledge of Pre-trained Text-Encoders for Continual Learning

### v1

@@ 1 Introduction
Alarge variety of algorithms were proposed to mitigate forgetting in continual learning, employing approaches such as rehearsal-based methods [1, 2, 3, 4], regularization-based techniques [5, 6, 7], and architecture-based solutions [8, 9, 10, 11, 12]. The emergence of Vision Transformers (ViTs) [13] has substantially enhanced the representation capabilities of pre-trained vision encoders for downstream tasks. A number of recent studies, including L2P [14] and DualPrompt [15], have leveraged the potential of pre-trained vision encoders in the context of continual learning, employing a prompt-based learning strategy. However, these works have exposed a notable limitation: performance degradation occurs when the pre-trained dataset exhibits a substantial semantic gap compared to the downstream data [16, 17].
Large scale pre-trained vision-language models, for instance ViT-BERT [18], CLIP [19] and UNITER [20], have emerged as powerful tools that combine computer vision and natural language processing, enabling machines to comprehend visual and textual information. These models can serve as strong foundation models for continual learning, as they were trained on massive datasets, allowing them to learn intricate patterns and relationships between images and language. As mentioned, the continual learning community has focused on applying foundational vision backbones [14, 15, 21], however, less effort has been invested in how to best exploit the rich semantic knowledge contained in the language encoders. A recent exception is Continual-CLIP [22] which shows impressive results in continual learning using both the vision and language encoder of the CLIP model. However, they keep the CLIP model frozen, thereby limiting the generalization to new tasks. In Fig. 1(a), we present an overview of several methods, each characterized by a distinct training strategy. The performance evaluations are shown in Fig. 1(b). The use of pre-trained vision encoders has resulted in incremental performance gains [^1] The methods L2P and DualPrompt obtain similar results without using any exemplars compared to existing state-of-the-art CL methods from scratch with exemplars (DER, DyTox++).. Interestingly, we will show that better exploitation of the knowledge contained in language encoders can lead to considerable performance improvements.
Method Language Vision Trainable
@@
L2P[14] x pre-trained yes
DualPrompt[15] x pre-trained yes
Continual-CLIP[22] pre-trained pre-trained no
CoOp[23] pre-trained pre-trained yes
@@
Figure 1: (a) Comparison of different models. (b) Performance comparison using different models across two datasets.
Establishing similarity between objects is a fundamental aspect of human perception and cognition, as highlighted by previous research [24]. pre-trained language models provide access to the semantic similarity between classes. In this paper, we propose two techniques to exploit the rich semantic information of pre-trained language models. Firstly, we exploit the intra-task semantic relationships between class labels. Instead of only using the available ground truth label to train on the newly arriving data, we replace the label with a soft-assignment towards all current task classes based on intra-task semantic relations. This ensures that each sample contributes to aligning the vision backbone with respect to multiple semantic classes. Secondly, we propose to exploit the inter-task semantic relationships between old and new classes to improve stability during continual learning. For example, current class data for ‘motor bikes’ is expected to assign part of its prediction to the related previous task class of ‘bikes’ thereby preventing forgetting of previous classes. These inter-task relationships, derived from the language encoder, are used in a semantically-guided knowledge distillation technique.
@@
  • The usage of pre-trained models in continual learning has focused on pre-trained vision encoders. We show that exploiting the semantic information of the data labels with pre-trained language encoders can greatly enhance the accuracy of continual learning.
  • We extend the existing image distillation method based only on visual information with a term that exploits the knowledge of the labels of the current data with respect to the classes of previous tasks (inter-task semantic similarity). This improvement, called SG-KD, is shown to significantly increase the efficiency of distillation (and reduce catastrophic forgetting). In addition, we show how to exploit the intra-task semantic similarity (called SG-RL).
  • Extensive experimental results show that our approach surpasses all methods by a substantial margin on several datasets under both many-shot and few-shot settings. Results on CIFAR100 under the 10-step setting show that the proposed method can greatly improve the accuracy after the last task with 11.4 points compared to the state-of-the-art.
@@ 2.1 Continual Learning
Continual learning research has witnessed significant advancements in recent years, resulting in the development of three main categories of approaches: rehearsal-based, regularization-based, and architecture-based. Rehearsal-based methods store previous task exemplars or generate fake exemplars using generative models to combat catastrophic forgetting. iCaRL [3] combines exemplar rehearsal with knowledge distillation, UCIR [2] addresses classification bias through normalization, and WA [4] ensures consistency in weight vector norms between old and new classes. Regularization-based methods impose constraints on network parameters, with strategies such as estimating parameter importance (EWC [5]) and distilling output from old and new models (LwF [6]) to retain information about previous tasks. SDC [25] proposed compensating the feature drift to prevent forgetting on some regularization-based continual learning methods. Architecture-based methods assign specified parameters to tasks or dynamically expand the network architecture. HAT [10] blocks and activates specific parts for old tasks, while DER [11] utilizes task-specific feature extractors and tackles parameter growth through pruning. Transformer-based continual learning paradigms, like DyTox [9], utilize shared self-attention layers and task-specific tokens to achieve task-specialized embeddings.  [26] proposed to apply Masked Autoencoders (MAEs) for continual learning and introduced a bilateral MAE framework that integrates learning from both image-level and embedding-level.  [21] proposed to adaptively assign distillation loss weights by evaluating the relevance of each patch to the current task.
@@ 2.2 Continual Learning with Foundation Models
A recent emerging trend in the field of continual learning is to combine pre-trained vision transformers with parameter-efficient fine-tuning techniques to continuously adapt the model to a stream of incoming downstream tasks. Specifically, these techniques involve prompt tuning [27, 28], prefix tuning [29], Adapter[30, 31], LoRA [32] , etc. The core of applying such a technique to continual learning is to construct additional learnable parameters or modules to instruct pre-trained representations and select appropriate prompts during inference time. L2P [14] optimizes the cosine similarity between query features and learnable keys, the most relevant prompts are selected and prepended to the token sequence both during training and inference. DualPrompt [15] further subdividing the prompts into general prompts sharing across all tasks and expert prompts picked and optimized in the same way as [14]. CODAPrompt [33] proposes to reweight prompts through input-conditioned weights, facilitating end-to-end optimization of the query-key mechanism across tasks. Distinct from the aforementioned prompt-based approaches, another line of work aims to construct classifiers through the extraction of robust feature representations via pretrained models. SLCA [17] fine-tuned the full pretrained model and recommended using different learning rates for the representation layers and the classifier to address the forgetting problem in representation layers, they further incorporated prototype replay for post-hoc alignment of the classification layers to reduce bias in the classifier. RanPAC [34] suggested employing a frozen random projection layer to project pretrained features into a high-dimensional space, thereby improving linear separability, they additionally utilize an online LDA classifier to eliminate correlations between categories. EASE [35] proposed to employ distinct adapters for each task to acquire task-specific features, and construct a unified classifier by synthesizing prototypes for old classes through semantic relevance.
@@ 3 Preliminary of CLIP model
The CLIP model has demonstrated remarkable continual learning performance without any fine-tuning [22]. To enhance its performance on new tasks and adapt to evolving data distributions over time, we thus propose to fine-tune the last block of the vision encoder. [^2] We fine-tune different parts of the vision encoder in Table VII. Our experiments indicate that fine-tuning the last block of the encoder results in the best performance. We aim to leverage the strong generalization ability of the CLIP model while also updating its representation to better adapt to incoming data.
@@ 4 Method
Figure 3: An illustration of two components. (a) The SG-RL module calculates the text embedding similarity between the current labels using a pre-trained text encoder. It then generates semantically-guided labels that are more informative. (b) The SG-KD module constructs an inter-task similarity matrix based on text embeddings of previous and current task class labels. The most similar categories of the current class from old classes are selected by computing row-wise maximums (pink cells). The similarity value is added to the previous model’s prediction logits at the most similar position, and the rest are subtracted from in each row. The enhanced logits are used for knowledge distillation between the current and previous models.
To overcome the challenges of catastrophic forgetting during the training of pre-trained models, we introduce a novel approach that integrates semantic information guidance into the process of continual knowledge learning, as shown in Fig. 2. Leveraging the power of well-trained text embeddings, our proposed approach facilitates efficient interaction within and across task labels, leading to improved performance from two crucial perspectives. Firstly, we focus on learning more informative representations for new data by leveraging intra-task semantic similarity, thereby enhancing model plasticity (SG-RL module). Secondly, we establish a relationship between old and new tasks by incorporating inter-task semantic similarity during the model distillation process, ensuring stability (SG-KD module). In the following sections, we introduce these two components and explain how they work.
@@ 4.1 Intra-task Semantically-Guided Representation Learning (SG-RL)
As shown in the yellow part of Fig. 2, assume we have $K$ classes contained in the current task $t$, their labels can be denoted as $C^{t}$=$\{c^{t}_{1},c^{t}_{2},\cdots,c^{t}_{K}\}$(for simplicity, we omit the superscript $t$ in the following descriptions). We encode the text embeddings of these labels by the pre-trained language part of CLIP to obtain the normalized text embeddings $\mathbf{T}^{t}=\{T_{1},T_{2},\cdots,T_{K}\}$, where $||T_{k}||=1$ and $k\in\{1,2,\cdots,K\}$. After obtaining the text embeddings of the current labels, we acquire the intra-task semantic similarity by computing the cosine similarity between each pair of text embeddings. This yields a text similarity matrix $\mathbf{S}^{t\leftrightarrow t}$ which represents the intra-task similarity between the $i$-th and $j$-th labels; it can be described as follows:
@@ 5.1 Experimental Settings
TABLE II: The results on CIFAR100 are averaged over three class orders (seeds) under 10, 20, and 50 steps, following the setting proposed in [9]. The bold parts represent the best results, underlined parts indicate the second best, and up arrows show the improvement of the best over the second best. Re-implemented results are marked with an asterisk (*) .
@@
Methods #Param. Avg Last #Param. Avg Last #Param. Avg Last
Linear probe - - 83.1 - - 83.1 - - 83.1
Joint 7.09 - 83.9 7.09 - 83.9 7.09 - 83.9
UCIR [2] 11.22 58.7 43.4 11.22 58.2 40.6 11.22 56.9 37.1
BiC [51] 11.22 68.8 53.5 11.22 66.5 47.0 11.22 62.1 41.0
WA [4] 11.22 69.5 53.8 11.22 67.3 47.3 11.22 64.3 42.1
PODNet [8] 11.22 58.0 41.1 11.22 54.0 35.0 11.22 51.2 33.0
DER w/o p [11] 112.27 75.4 65.2 224.55 74.1 62.5 561.39 72.4 59.1
DER [11] - 74.6 64.4 - 74.0 62.6 - 72.1 59.8
DyTox++ [9] 10.73 77.0 67.5 10.74 76.8 64.3 10.77 75.5 59.5
@@
CoOp[23] 7.91 76.2 68.1 7.91 77.0 67.6 7.91 78.3 66.3
CoCoOp[38] 7.13 75.1 65.9 7.13 75.1 63.1 7.13 74.8 61.1
L2P[14] 7.38 76.3 65.9 7.38 75.2 66.2 7.38 76.5 64.6
DualPrompt[15] 7.42 71.3 65.1 7.42 66.2 65.4 7.42 70.1 66.6
Ours 7.09 86.6 ($\uparrow$9.6) 80.1 ($\uparrow$11.4) 7.09 86.0 ($\uparrow$9.0) 78.2 ($\uparrow$9.5) 7.09 81.8 ($\uparrow$3.5) 74.1 ($\uparrow$5.4)
TABLE III: We compare the results obtained on the imagenet_subset dataset (100 classes) and the ImageNet full dataset (1000 classes), where the training session is split into 10 steps, with each step containing an equal number of classes for each task.
@@
#Param. top-1 top-5 #Param. top-1 top-5
Avg Last Avg Last Avg Last Avg Last
Linear probe - - 83.9 - 97.6 - - 80.2 - 94.1
Joint 7.09 - 86.1 - 98.3 7.09 - 81.1 - 96.5
DER w/o p[11] 112.27 77.2 66.7 93.2 87.5 116.89 68.8 60.2 88.2 82.9
DER[11] - 76.1 66.1 92.8 88.4 - 66.7 58.6 87.1 81.9
DyTox++[9] 11.01 80.8 72.5 94.4 90.1 - - - - -
@@
CoOp[23] 7.91 80.9 70.2 97.5 94.5 15.28 75.2 66.7 95.9 91.9
CoCoOp[38] 7.13 81.8 71.0 98.5 96.1 7.13 64.6 56.0 92.3 88.7
L2P[14] 7.38 82.2 72.7 98.5 96.4 8.07 71.9 63.3 95.0 90.4
DualPrompt[15] 7.42 79.6 68.4 97.7 95.0 8.11 70.7 64.0 94.4 92.1
Ours 7.09 89.8 ($\uparrow$7.6) 83.1 ($\uparrow$7.9) 99.1 ($\uparrow$0.6) 98.4 ($\uparrow$1.5) 7.09 83.4 ($\uparrow$8.2) 75.1 ($\uparrow$6.5) 97.4 ($\uparrow$1.5) 95.2 ($\uparrow$3.1)
@@
Our experimental evaluation begins by conducting experiments on three datasets that are commonly used for continual learning scenarios: CIFAR100 [42], imagenet_subset, and ImageNet1000  [43]. For CIFAR100, we adopt the incremental phase splitting into 10, 20, and 50 steps, following the approach of DyTox [9]. In each split, we further assess the performance of the proposed method under three different class orders. Regarding imagenet_subset and ImageNet1000, we split the incremental phase into 10 steps, with 10/100 new classes added at each incremental step. We further evaluate the performance on eight fine-grained datasets. We also evaluate for few-shot continual learning on miniImagenet, and CUB-200-2011 following [52, 53]. For miniImagenet, we split the 100 classes into 60 base classes and 40 classes across 8 sessions with 5 new classes per session. CUB-200-2011 is divided into 100 base classes and 100 classes across 10 sessions with 10 new classes per session. Each incremental session for all three datasets consists of 5 training samples per new class.
@@
All experiments were conducted with a vision backbone of ViT-B/16 version CLIP model. We train every task for 10 epochs except for CUB and Aircrafts with 20 epochs. We adopt SGD in all experiments as the optimizer with an initial learning rate of 0.01(0.001 for few-shot setting), weight decay is 2e-4 and 0.9 for momentum. Batch size is 256 for all experiments. We set $\alpha$ to 13 in Eq. 6, $\beta$ is set to 100 in Eq. 8 following [22] and $\tau$ to 0.1 in Eq. 14 for all datasets. The trade-off parameters, $\lambda_{1}$ and $\lambda_{2}$, are set to 0.5 and 0.1 across all datasets (the hyper-parameters sensitive experiments can be found in Section 5.4.5 ). We save 20 exemplars for each old class by herding algorithm in all experiments except the ablation on different exemplar sizes (in Section 5.4.4), following the setting in [9, 3, 11]. Under the few-shot setting, none exemplar is saved following the setting in [52, 53]. To ensure fairness, methods trained with a pre-trained model are initialized with CLIP [^4] Since the exact prompt and specific class names employed by CLIP [19] for each dataset are unknown. We tried several prompts and finally for each dataset we utilize a specific prompt so that the results of our own implementation are comparable or identical to those provided by CLIP [19].. For implementing ‘Linear Probe’, we directly add a fully connected layer after the linear projection, we train 100 epochs for each dataset with an initial learning rate of 0.1 and decays by 0.1 every 45 epochs. For ‘Joint’, to ensure fairness in comparison, we fine-tune the last block of the network for joint training, all training samples are acquired concurrently in the same session. We train 100 epochs for each dataset with an initial learning rate of 0.01 and decays by 0.1 every 45 epochs.
For all experiments in the paper, we applied RandomResizedCrop, RandomHorizontalFlip and normalization for data augmentation. The random seed was set to 1993 except when spliting CIFAR100 with three different class orders as shown in Table 2. We use randomly generated integers ranging from 1 to 5000 for the other two class orders.
Test-time evaluation details: For each test sample $x_{\text{test}}$, the image is passed through the image encoder, resulting in its image feature $I_{\text{test}}$. Meanwhile, we collect all previously seen categories as prompts and feed them to the text encoder to obtain the text features $T=\left\{T^{i};1\leq i\leq k\right\}$, where $k$ is the number of seen classes. Then we compute the cosine similarity between image feature and text features:
$s_{i}=\frac{I_{test}\cdot T_{i}}{\left\|I_{test}\left\|\cdot\right\|T_{i}\right\|}$ (16)
where $1\leq i\leq k$, the prediction score for $x_{test}$ can then be computed as:
$p=\frac{\exp(s_{i})}{{\textstyle\sum_{i=1}^{k}}\exp(s_{i})}$ (17)
The class with the highest predicted score is considered as the predication.

### v2

@@ 1 Introduction
Continual learning (CL) aims to enable models to learn sequentially from a stream of tasks without catastrophically forgetting previously acquired knowledge. A large variety of algorithms were proposed to mitigate forgetting, employing approaches such as rehearsal [1, 2, 3, 4], regularization [5, 6, 7], and architectural modifications [8, 9, 10, 11, 12]. The emergence of Vision Transformers (ViTs) [13] has substantially enhanced the representation capabilities of pre-trained vision encoders for downstream tasks. A number of recent studies, including L2P [14] and DualPrompt [15], have leveraged the potential of pre-trained vision encoders in the context of continual learning, employing a prompt-based learning strategy. However, these works have exposed a notable limitation: performance degradation occurs when the pre-trained dataset exhibits a substantial semantic gap compared to the downstream data [16, 17].
Large scale pre-trained vision-language models (VLMs), for instance ViT-BERT [18], CLIP [19] and UNITER [20], have emerged as powerful tools that combine computer vision and natural language processing, enabling machines to comprehend visual and textual information [21, 22]. These models can serve as strong foundation models for continual learning, as they were trained on massive datasets, allowing them to learn intricate patterns and relationships between images and language. As mentioned, the continual learning community has focused on applying foundational vision backbones [14, 15, 23, 24, 25]. However, how to best exploit the rich semantic knowledge embedded within the language encoders of VLMs is an active and evolving area of investigation. Early efforts like Continual-CLIP [26] demonstrated the potential by utilizing both vision and language encoders of CLIP, albeit by keeping the entire model frozen, which can limit adaptability to new tasks. Language guidance from pre-trained language encoders has been introduced into prompt-based continual learning methods, such as LGCL [27]. This approach aligns the task selection mechanism in the prompt pool with the overall language description of the task and aligning the visual features with the language description of specific classes.
Method Language Vision Learnable Backbone
@@
CODA-Prompt[23] x pre-trained no
EASE [25] x pre-trained no
ConvPrompt [24] x pre-trained no
RAPF [28] pre-trained pre-trained no
CLAP4CLIP [29] pre-trained pre-trained no
@@
Figure 1: (a) Comparison of different models. (b) Performance comparison using different models across three datasets.
Establishing similarity between objects is a fundamental aspect of human perception and cognition, as highlighted by previous research [30]. Pre-trained text encoders provide access to the semantic similarity between classes. In this paper, we propose two techniques to exploit the rich semantic information of pre-trained text encoders. Firstly, we exploit the intra-task semantic relationships between class labels. Instead of only using the available ground truth label to train on the newly arriving data, we replace the label with a soft-assignment towards all current task classes based on intra-task semantic relations. This ensures that each sample contributes to aligning the vision backbone with respect to multiple semantic classes. Secondly, we propose to exploit the inter-task semantic relationships between old and new classes to improve stability during continual learning. For example, current class data for ‘motor bikes’ is expected to assign part of its prediction to the related previous task class of ‘bikes’ thereby preventing forgetting of this class. These inter-task relationships, derived from the language encoder, are used in a semantically-guided knowledge distillation technique.
In Fig. 1(a), we present an overview of several methods, each characterized by a distinct training strategy. The performance evaluations are shown in Fig. 1(b). We observe that the incorporation of pre-trained vision-language models leads to performance improvements on ImageNet-R, which poses a large semantic gap from CLIP’s pre-training distribution. This suggests that the rich semantic priors from language supervision enhance the model’s generalization to out-of-distribution categories. However, methods with frozen backbones, such as CLAP4CLIP [29], exhibit clear limitations when facing fine-grained and domain-shifted datasets like Aircraft, likely due to their inability to adapt feature representations to new tasks. In contrast, our approach effectively combines the benefits of vision-language pre-training with end-to-end adaptability, achieving superior performance across all datasets and outperforming both vision-only and frozen vision-language baselines under challenging class-incremental learning scenarios.
@@
  • We demonstrate that effectively leveraging the semantic knowledge from pre-trained language encoders, in conjunction with vision encoders, significantly enhances performance in class-incremental learning, moving beyond the predominant focus on vision-only pre-training or less direct uses of language information.
  • We introduce the SG-RL module, which leverages intra-task semantic similarity between class labels to replace conventional one-hot label supervision. By aligning features with multiple semantically related class embeddings rather than treating classes as mutually independent, this module encourages the model to capture richer semantic structure within each task, leading to more discriminative and transferable representations.
  • We extend the existing image distillation method based only on visual information with a term that exploits knowledge of the labels of the current data with respect to the classes of previous tasks (inter-task semantic similarity). This improvement, called SG-KD, significantly increases the efficiency of distillation (and reduces catastrophic forgetting).
  • Extensive experimental validation across a wide range of benchmarks—including the standard CIFAR100, large-scale datasets like ImageNet-subset and ImageNet-1000, eight diverse fine-grained datasets and challenging few-shot learning scenarios —confirms that our approach consistently delivers state-of-the-art or highly competitive performance.
@@ 2.1 Continual Learning
Continual learning research has witnessed significant advancements in recent years, early approaches [3, 5, 6, 31, 32, 33, 34, 35] can be broadly divided into three categories: rehearsal-based, regularization-based, and architecture-based. Rehearsal-based methods store previous task exemplars or generate fake exemplars using generative models to combat catastrophic forgetting. iCaRL [3] combined exemplar rehearsal with knowledge distillation, UCIR [2] addressed classification bias through normalization, and WA [4] ensured consistency in weight vector norms between old and new classes. Regularization-based methods impose constraints on network parameters, with strategies such as estimating parameter importance (EWC [5]) and distilling output [36, 37] from old and new models (LwF [6]) to retain information about previous tasks. SDC [31] proposed compensating the feature drift to prevent forgetting. Architecture-based methods assign specified parameters to tasks or dynamically expand the network architecture. HAT [10] blocked and activated specific parts for old tasks, while DER [11] utilized task-specific feature extractors and tackles parameter growth through pruning. Transformer-based continual learning paradigms, like DyTox [9], utilized shared self-attention layers and task-specific tokens to achieve task-specialized embeddings. CMAE [38] proposed to apply Masked Autoencoders (MAEs) for continual learning and introduced a bilateral MAE framework that integrates learning from both image-level and embedding-level. FKSR [39] proposed to adaptively assign distillation loss weights by evaluating the relevance of each patch to the current task.
@@ 2.2 Continual Learning with Foundation Models
A recent emerging trend in the field of continual learning is to combine pre-trained vision transformers with parameter-efficient fine-tuning techniques to continuously adapt the model to a stream of incoming downstream tasks. Specifically, these techniques involve prompt tuning [40, 41], prefix tuning [42], Adapter[43, 44], LoRA [45] , etc. The core of applying such a technique to continual learning is to construct additional learnable parameters or modules to instruct pre-trained representations and select appropriate prompts during inference time. L2P [14] optimized the cosine similarity between query features and learnable keys, the most relevant prompts are selected and prepended to the token sequence both during training and inference. DualPrompt [15] further subdivided the prompts into general prompts sharing across all tasks and expert prompts picked and optimized in the same way as [14]. CODA-Prompt [23] proposed to reweight prompts through input-conditioned weights, facilitating end-to-end optimization of the query-key mechanism across tasks. ConvPromt [24] proposed to generate layer-wise prompts from shared embeddings and integrated Large Language Model-driven task similarity analysis to dynamically allocate prompt resources. Distinct from the aforementioned prompt-based approaches, another line of work aims to construct classifiers through the extraction of robust feature representations via pre-trained models. SLCA [17] fine-tuned the full pre-trained model and recommended using different learning rates for the representation layers and the classifier to address the forgetting problem in representation layers, they further incorporated prototype replay for post-hoc alignment of the classification layers to reduce bias in the classifier. RanPAC [46] suggested employing a frozen random projection layer to project pre-trained features into a high-dimensional space, thereby improving linear separability, they additionally utilize an online LDA classifier to eliminate correlations between categories. EASE [25] proposed to employ distinct adapters for each task to acquire task-specific features, and construct a unified classifier by synthesizing prototypes for old classes through semantic relevance. Considering the feature generalizability of large-scale vision-language pre-trained models, a series of recent works has emerged that either specifically design for or leverage these types of models [47, 48, 49, 50, 28, 29, 51]. SPU [50] introduced a selective parameter update strategy to mitigate generic knowledge forgetting in foundation models, by localizing to key layers and fine-tuning only a small portion of task-relevant parameters. CLAP4CLIP [29] addressed uncertainties in multi-modal interactions by probabilistically modeling the visual-guided text feature space and incorporating task-specific probabilistic adapter modules. Furthermore, it leverages CLIP’s rich pre-trained knowledge for weight initialization and distribution regularization to effectively alleviate catastrophic forgetting. RAPF [28] proposed to adaptively adjusting affected old class representations using textual features, combined with a decomposed parameter fusion technique to stabilize the adapter. PROOF [51] proposed to learn new knowledge by training task-specific expandable projection layers and freezing old ones to prevent forgetting, while also designing a cross-modal fusion module to better integrate visual and textual information.
@@ 3 Preliminary of CLIP model
The CLIP model has demonstrated remarkable continual learning performance without any fine-tuning [26]. To enhance its performance on new tasks and adapt to evolving data distributions over time, we thus propose to fine-tune the last block of the vision encoder. [^1] We fine-tune different parts of the vision encoder in Table VIII. Our experiments indicate that fine-tuning the last block of the encoder results in the best performance. We aim to leverage the strong generalization ability of the CLIP model while also updating its representation to better adapt to incoming data.
@@ 4 Method
Figure 3: An illustration of two components. (a) The SG-RL module calculates the text embedding similarity between the current labels using a pre-trained text encoder. It then generates semantically-guided labels that are more informative. (b) The SG-KD module constructs an inter-task similarity matrix based on text embeddings of previous and current task class labels. By introducing semantic relationships between old and new classes without requiring any visual information, it enables the model to capture inter-task semantic connections and perform knowledge transfer, thereby enhancing stability.
To overcome the challenges of catastrophic forgetting during the training of pre-trained models, we introduce a novel approach that integrates semantic information guidance into the process of continual knowledge learning, as shown in Fig. 2. Leveraging the power of well-trained text embeddings, our proposed approach facilitates efficient interaction within and across task labels, leading to improved performance from two crucial perspectives. Firstly, we focus on learning more informative representations for new data by leveraging intra-task semantic similarity, thereby enhancing model plasticity (SG-RL module). Secondly, we establish a relationship between old and new tasks by incorporating inter-task semantic similarity during the model distillation process, ensuring stability (SG-KD module). Similar to recent methods [29, 51, 50], we also use a small set of exemplars in the form of old task images and text labels. In the following sections, we introduce these two components and explain how they work.
@@ 4.1 Intra-task Semantically-Guided Representation Learning (SG-RL)
As shown in the yellow part of Fig. 2, assume we have $K$ classes contained in the current task $t$ (including exemplar classes, if applicable), their labels can be denoted as $C^{t}$=$\{c^{t}_{1},c^{t}_{2},\cdots,c^{t}_{K}\}$(for simplicity, we omit the superscript $t$ in the following descriptions). We encode the text embeddings of these labels by the pre-trained language part of CLIP to obtain the normalized text embeddings $\mathbf{T}^{t}=\{T_{1},T_{2},\cdots,T_{K}\}$, where $||T_{k}||=1$ and $k\in\{1,2,\cdots,K\}$. After obtaining the text embeddings of the current labels, we acquire the intra-task semantic similarity by computing the cosine similarity between each pair of text embeddings. This yields a text similarity matrix $\mathbf{S}^{t\leftrightarrow t}$ which represents the intra-task similarity between the $i$-th and $j$-th labels; it can be described as follows:
@@ 5.1 Experimental Settings
ImageNet-A  [67] 200 5981 1519
ImageNet-R  [68] 200 24000 6000
VTAB  [69] 50 1796 8619
TABLE II: The results on CIFAR100 under 10, 20, and 50 steps. The bold parts represent the best results, underlined parts indicate the second best. PTM-based methods were re-implemented with the same CLIP ViT-B/16 weights for fair comparison. We denote if the methods could be used for zero-shot classification (ZS) which is true (✓) for CLIP-based methods and false (✗) for methods using only vision backbones. We also denote the methods with learnable backbone (LB) by (✓) and the prompt- or adapter-based methods having frozen backbone by (✗).
@@
Methods ZS LB Avg Acc. Last Acc. Avg Acc. Last Acc. Avg Acc. Last Acc.
Linear probe - - - 83.1 - 83.1 - 83.1
Joint - - - 89.3 - 89.3 - 89.3
DER w/o p[11] 75.4 65.2 74.1 62.5 72.4 59.1
DER[11] 74.6 64.4 74.0 62.6 72.1 59.8
DyTox++[9] 77.0 67.5 76.8 64.3 75.5 59.5
@@
DualPrompt [15] 76.8 67.4 72.9 66.4 73.0 70.3
RanPAC [46] 88.6 82.0 88.9 82.0 89.0 81.8
CODA-Prompt [23] 78.1 67.8 71.4 60.3 47.8 36.0
ConvPrompt [24] 80.1 70.3 73.7 62.5 53.2 40.1
EASE [25] 82.4 73.4 79.8 67.9 53.8 32.7
[…]
CoOp [55] 74.9 66.6 75.1 65.9 79.2 69.9
CoCoOp [54] 71.6 64.1 77.3 67.5 80.5 71.5
SPU [50] 87.6 80.0 87.4 80.5 86.9 78.8
CLAP4CLIP [29] 86.5 78.6 86.0 77.4 82.6 75.0
RAPF [28] 86.4 79.6 86.0 78.0 81.3 69.9
PROOF [51] 83.0 73.8 83.4 73.6 83.5 72.7
Ours 89.6 83.3 89.6 82.8 89.5 82.2
TABLE III: We compare the results on the ImageNet_subset dataset (100 classes) and the ImageNet1000 full dataset (1000 classes), where the training session is split into 10 steps, with each step containing an equal number of classes for each task.
@@
top-1 top-5 top-1 top-5
ZS LB Avg Last Avg Last Avg Last Avg Last
Linear probe - - - 83.9 - 97.6 - 80.2 - 94.1
Joint - - - 86.1 - 98.3 - 81.1 - 96.5
DER w/o p[11] 77.2 66.7 93.2 87.5 68.8 60.2 88.2 82.9
DER[11] 76.1 66.1 92.8 88.4 66.7 58.6 87.1 81.9
DyTox++[9] 80.8 72.5 94.4 90.1 73.2 64.6 91.1 87.1
@@
DualPrompt [15] 81.7 71.3 98.5 96.9 73.4 65.6 95.6 93.2
RanPAC [46] 90.7 83.7 98.3 96.8 81.8 73.9 96.2 93.7
CODA-Prompt [23] 81.1 66.8 97.8 95.3 81.7 73.6 97.0 94.9
ConvPrompt [24] 73.3 65.0 94.1 88.1 76.2 70.8 94.5 92.2
EASE [25] 83.6 71.0 98.5 96.6 79.1 70.3 96.3 93.7
[…]
CoOp [55] 82.2 72.0 98.8 97.6 75.1 67.5 96.3 89.7
CoCoOp [54] 80.1 69.1 97.9 95.4 69.2 58.9 89.4 82.3
SPU [50] 86.5 77.6 97.6 95.2 78.7 70.7 94.1 91.0
CLAP4CLIP [29] 87.0 79.1 96.6 93.8 81.5 74.0 94.7 91.5
RAPF [28] 88.3 80.2 98.9 97.5 79.9 67.3 95.1 89.5
PROOF [51] 82.3 65.6 98.1 95.4 70.1 56.5 92.9 87.2
Ours 89.8 83.1 99.1 98.4 83.4 75.1 97.4 95.2
@@
Our experimental evaluation begins by conducting experiments on three datasets that are commonly used for continual learning scenarios: CIFAR100 [58], ImageNet_subset, and ImageNet1000  [59]. For CIFAR100, we adopt the incremental phase splitting into 10, 20, and 50 steps, following the approach of DyTox [9]. Regarding ImageNet_subset and ImageNet1000, we split the incremental phase into 10 steps, with 10/100 new classes added at each incremental step. We further evaluate the performance on eight fine-grained datasets. We also evaluate for few-shot continual learning on miniImageNet, and CUB-200-2011 following [70, 71]. For miniImageNet, we split the 100 classes into 60 base classes and 40 classes across 8 sessions with 5 new classes per session. CUB-200-2011 is divided into 100 base classes and 100 classes across 10 sessions with 10 new classes per session. Each incremental session for all three datasets consists of 5 training samples per new class.
@@
We train every task for 10 epochs except for CUB and Aircrafts with 20 epochs. We adopt SGD in all experiments as the optimizer with an initial learning rate of 0.01(0.001 for few-shot setting), weight decay is 2e-4 and 0.9 for momentum. Batch size is 256 for all experiments. We set $\alpha$ to 13 in Eq. 6, $\beta$ is set to 100 in Eq. 8 following [26] and $\tau$ to 0.1 in Eq. 14 for all datasets. The trade-off parameters, $\lambda_{1}$ and $\lambda_{2}$, are set to 0.5 and 0.1 across all datasets (the hyper-parameters sensitive experiments can be found in Section 5.5.5 ). We save 20 exemplars for each old class by herding algorithm in all experiments except the ablation on different exemplar sizes (in Section 5.5.4), following the setting in [9, 3, 11]. Under the few-shot setting, none exemplar is saved following the setting in [70, 71]. For all PTM based methods, we re-implemented with the same pre-trained ViT-B/16 CLIP [19] weights for fair comparison.333 For Continual-CLIP [26], since the exact prompt and specific class names employed by CLIP [19] for each dataset are unknown. We tried several prompts and finally for each dataset we utilize a specific prompt so that the results of our own implementation are comparable or identical to those provided by CLIP [19].. For implementing ‘Linear Probe’, we directly add a fully connected layer after the linear projection, we train 100 epochs for each dataset with an initial learning rate of 0.1 and decays by 0.1 every 45 epochs. For ‘Joint’, all training samples are acquired concurrently in the same session. We train 100 epochs for each dataset with an initial learning rate of 0.01 and decays by 0.1 every 45 epochs.
For CIFAR100, we refined this general configuration to ensure fair comparison with recent highly-optimized methods [50], given its unique low-resolution nature. Specifically, for CIFAR100 only, we activated all attention projection matrices in residual blocks from shallow to deep layers to better capture low-level features, we set the learning rate to 5e-6 and adopted the AdamW optimizer.
For all re-implementation details of compared methods, please refer to Appendix D.
5.1.3 Evaluation metrics
Following the evaluation protocol of previous works [9, 14, 17], we report the average incremental accuracy ‘Avg’ and the final accuracy ‘Last’. For a sequence of tasks denoted as $\mathcal{T}=\{1,2,\dots,n\}$, let $A_{t}$ represent the mean accuracy on all learned categories after learning task $t$, then ‘Avg’ can be denoted as $Avg=\frac{1}{n}{\textstyle\sum_{1}^{n}}A_{t}$, and the final accuracy ‘Last’ is $A_{n}$ which represent the average accuracy of all categories after learning the last task $n$. Unless explicitly stated, all experiments and ablation studies herein will report only the final accuracy ‘Last’.